# Fast Streaming Transducer ASR Prototyping via Knowledge Distillation with Whisper

### v1

@@ Abstract
Iuliia Thorbecke [^1] Equal contribution. Order is determined by a coin flip.1,2 Juan Zuluaga-Gomez [^1] Equal contribution. Order is determined by a coin flip.1,3 Esaú Villatoro-Tello1 Shashi Kumar1,3 Pradeep Rangappa1 Sergio Burdisso1 Petr Motlicek1,4 Karthik Pandia5 Aravind Ganapathiraju5 1 Idiap Research Institute 2 University of Zurich 3 EPFL 4 Brno University of Technology 5 Uniphore iuliia.thorbecke@uzh.ch juan.zuluaga@eu4m.eu
@@ 1 Introduction
Figure 1: Proposed framework for efficient and fast streaming ASR prototyping with pseudo-labeled data. Transducer model are further improved via shallow fusion of n-gram LMs and contextual biasing of target named entities.
@@
The Transformer-Transducer architecture (Yeh et al., 2019) is widely exploited for industrial uses that require streaming decoding because the transducer decoder naturally supports streaming (Li et al., 2021, 2020). However, the transducer used to be harder to train compared to AED and CTC, thus, it was less explored in the community, until it was shown to achieve a performance as close as AED models (Sainath et al., 2020). The transducer models consist of an encoder, predictor and joint networks. Using a Transformer (Vaswani et al., 2017) encoder leads to a Transformer-Transducer (TT) architecture (Battenberg et al., 2017; Yeh et al., 2019; Zhang et al., 2020). When trained from scratch, the TT models require sufficient amounts of supervised datasets in the target language and domain (Noroozi et al., 2023; Li et al., 2021). At the same time, fine-tuning a large pre-trained model, even when using a transducer decoder, would not allow streaming decoding.
@@ 2 Related Work
Semi-supervised learning (Higuchi et al., 2021), pseudo labeling (Zavaliagkos and Colthurst, 1998), and weakly supervised learning (Radford et al., 2022) are a family of methods aiming to partly alleviate the burden of lack of labeled data for supervised ASR training. These approaches have shown promising word error rate (WER) improvement in multiple settings and languages. In practice, a teacher model $g$ is trained on an audio-text paired corpus $D_{l}=\{X_{i},Y_{i}\}$. Then, it is used to pseudo label a much larger unlabeled only audio corpus, $D_{pl}=\{X_{i},Y^{*}_{i}\}$. Afterward, usually a smaller model (Barrault et al., 2023) can use $D_{l}$ and $D_{pl}$ for supervised training or fine-tuning (Hsu et al., 2021). However, PLs are often noisy and bounded by $g$ model quality, whereas their use might result in suboptimal final performance in the models. This can be solved by either filtering out the nosiest samples or increasing $g$ model size to improve their quality. [^2] We assume that a larger model, trained under the same conditions and with increased data, will attain lower WERs. Several approaches to improve the PL quality include improving loss functions (Zhu et al., 2023; Gao et al., 2023), pairing online and offline models at training time (Higuchi et al., 2021), and continuous single-language  (Likhomanenko et al., 2022; Berrebbi et al., 2022) and multilingual pseudo-labeling setting (Lugosch et al., 2022).
@@ 3.2 Transformer-Transducer Training
The peak learning rate is $lr=5.0e^{-2}$ and we train each TT for 30 epochs on a single RTX 3090 GPU with only PLs. [^8] We also run an experiment valuable for the industrial domain. It includes a thorough analysis of PLs quality for the call-center domain, see Appendix D. Training takes between 1 and 2 days.
@@ 3.4.1 CommonVoice Database
The CommonVoice dataset comprises several thousand hours of audio in more than 100 languages (Ardila et al., 2020). For experimentation, we select six languages from CommonVoice-v11 (Ardila et al., 2020)  [^16] CommonVoice-v11: cv-corpus-11.0-2022-09-21 which have sufficient data for training ASR and language models: Catalan (CA), English (EN), German (DE), French (FR), Spanish (ES), and Italian (IT). We use the official train sets and report WERs on the official test sets. See Table 1 for further statistics.
@@ 4.1 Performance on models with PL of different quality
Low-latency streaming decoding
5 Conclusions
In this work, we propose a framework to meet the challenge of training streaming ASR systems with few-to-none supervised data by leveraging PLs from foundational speech models. We conduct a thorough examination of the efficacy of PL-based TT models across various dimensions, including offline and chunk-wise decoding for streaming applications, and the influence of FSM size on the TT model’s WERs. We introduce robust heuristics to filter out unreliable and hallucinated PLs. Our findings reveal that TT models can be effectively trained from scratch on noisy PLs. We managed to further improve the performance of models trained with weak pseudo-labels (generated by Whisper-tiny, -base, and -small) by adding regularization with different amounts of supervised data. Additionally, we prove that decoding with the shallow fusion of external n-gram LM and automatically generated named entities always improves the performance of models independent of the quality of pseudo-labels.
Limitations
Ethical Considerations
Appendix A Streaming decoding configurations
  • Decode chunk size = 320ms with left context of 2560ms, 5120ms and full;
  • decode chunk size = 640ms with left context of 2560ms, 5120ms and full;
Appendix B Extended results for models trained on PL data
Offline models evaluation
Table 3: WERs for six CommonVoice languages. The Zipformer offline models are trained with pseudo-labeled data from different Whisper models. We also report WERs when a small amount of supervised data is added during training, denoted as "sup. [100h]". Note that the transducer models are trained from scratch in $\sim$1 day GPU time.
@@ Appendix B Extended results for models trained on PL data
Streaming models evaluation

### v2

@@ Abstract
Iuliia Thorbecke [^1] Equal contribution. Order is determined by a coin flip.1,2 Juan Zuluaga-Gomez [^1] Equal contribution. Order is determined by a coin flip.1,3 Esaú Villatoro-Tello1 Shashi Kumar1,3 Pradeep Rangappa1 Sergio Burdisso1 Petr Motlicek1,4 Karthik Pandia5 Aravind Ganapathiraju5 1 Idiap Research Institute, Switzerland; 2 University of Zurich, Switzerland; 3 EPFL, Switzerland; 4 Brno University of Technology, Czech Republic; 5 Uniphore, India iuliia.nigmatulina@idiap.ch juan.zuluaga@eu4m.eu
@@ 1 Introduction
Figure 1: Proposed framework for efficient and fast streaming ASR prototyping with pseudo-labeled data. Transducer models are further improved via shallow fusion of n-gram LMs and contextual biasing of target named entities.
@@
The Transformer-Transducer architecture (Yeh et al., 2019) is widely exploited for industrial uses that require streaming decoding because the transducer decoder naturally supports streaming (Li et al., 2021, 2020). However, the transducer used to be harder to train compared to AED and CTC, thus, it was less explored in the community, until it was shown to achieve a performance as close as AED models (Sainath et al., 2020). The transducer models consist of an encoder, predictor and joint networks. Using a Transformer (Vaswani et al., 2017) encoder leads to a Transformer-Transducer (TT) architecture (Battenberg et al., 2017; Yeh et al., 2019; Zhang et al., 2020a). When trained from scratch, the TT models require sufficient amounts of supervised datasets in the target language and domain (Noroozi et al., 2023; Li et al., 2021). At the same time, fine-tuning a large pre-trained model, even when using a transducer decoder, would not allow streaming decoding.
@@ 2 Related Work
Semi-supervised learning (Zhang et al., 2020b; Park et al., 2020; Higuchi et al., 2021), pseudo labeling (Zavaliagkos and Colthurst, 1998; Likhomanenko et al., 2020; Hwang et al., 2022), and weakly supervised learning (Radford et al., 2022) are a family of methods aiming to partly alleviate the burden of lack of labeled data for supervised ASR training. These methods have shown promising word error rate (WER) improvement in multiple settings and languages. In practice, a teacher model is trained on an audio-text paired corpus $D_{l}=\{X_{i},Y_{i}\}$. Then, it is used to pseudo label a much larger unlabeled only audio corpus, $D_{pl}=\{X_{i},Y^{*}_{i}\}$. Afterward, usually a smaller model (Barrault et al., 2023) can use $D_{l}$ and $D_{pl}$ for supervised training or fine-tuning (Hsu et al., 2021).
The main difference between most of the previous studies and our approach proposed in the paper is that typically iterative training is used Zhang et al. (2020b); Park et al. (2020); Hwang et al. (2022). A multi-stage strategy combining self- and semi-supervised learning eventually results in strong pseudo-labels. In the present paper, we focus on the performance that can be achieved “out-of-the-box” by using already available FSM and training transducer models from scratch and only once. Thus, this approach allows for minimizing the overall computational cost, i.e., one-stage training and applying improvement methods, such as decoding with shallow fusion. We aim to reveal the potential of the FSM in pseudo-label generation and demonstrate what performance can be reached with minimal training efforts.
PLs, however, are often noisy and bounded by the quality of a teacher model, whereas their use might result in suboptimal final performance in the models. This can be solved by either filtering out the nosiest samples or increasing the teacher model size to improve their quality. [^2] We assume that a larger model, trained under the same conditions and with increased data, will attain lower WERs. Several approaches to improve the PL quality include improving loss functions (Zhu et al., 2023; Gao et al., 2023), pairing online and offline models at training time (Higuchi et al., 2021), and continuous single-language  (Likhomanenko et al., 2022; Berrebbi et al., 2022) and multilingual pseudo-labeling setting (Lugosch et al., 2022).
@@ 3.2 Transformer-Transducer Training
We use an effective batch size of 600s with a gradient accumulation of 1, the peak learning rate is $lr=5.0e^{-2}$ and we train each TT for 30 epochs on a single RTX 3090 GPU with only PLs. [^8] We also run an experiment valuable for the industrial domain. It includes a thorough analysis of PLs quality for the call-center domain, see Appendix D. Training takes between 1 and 2 days. During the decoding, we use a beam size of 4.
@@ 3.4.1 CommonVoice Database
The CommonVoice dataset comprises several thousand hours of audio in more than 100 languages (Ardila et al., 2020). To the best of our knowledge, [^16] According to the discussions in the official OpenAI-Whisper GitHub repository: https://github.com/openai/whisper/discussions/349, https://github.com/openai/whisper/discussions/2305. the CommonVoice data was not used for training Whisper model and can be used for zero-shot evaluation. In our case, it is an important point, as using unseen data for generating PLs provides a more realistic estimation of the proposed approach performance. [^17] Although officially the CommonVoice data is not included in the training data for Whisper, we realize the possibility of some CommonVoice data still being seen by the model through other sources and in a small amount. For experimentation, we select six languages from CommonVoice-v11 (Ardila et al., 2020)  [^18] CommonVoice-v11: cv-corpus-11.0-2022-09-21 which have sufficient data for training ASR and language models: Catalan (CA), English (EN), German (DE), French (FR), Spanish (ES), and Italian (IT). We use the official train sets and report WERs on the official test sets. See Table 1 for further statistics.
@@ 4.1 Performance on models with PL of different quality
Table 3: WERs for six CommonVoice languages. The Zipformer offline models are trained with pseudo-labeled data from different Whisper models. We also report WERs when a small amount of supervised data is added during training, denoted as ‘‘sup. [100h]’’. Note that the transducer models are trained from scratch in $\sim$1 day GPU time.
@@
Low-latency streaming decoding
5 Conclusions
In this work, we propose a framework to meet the challenge of training streaming ASR systems with few-to-none supervised data by leveraging PLs from foundational speech models. We conduct a thorough examination of the efficacy of PL-based TT models across various dimensions, including offline and chunk-wise decoding for streaming applications, and the influence of FSM size on the TT model’s WERs. We introduce robust heuristics to filter out unreliable and hallucinated PLs. Our findings reveal that TT models can be effectively trained from scratch on noisy PLs. We managed to further improve the performance of models trained with weak pseudo-labels (generated by Whisper-tiny, -base, and -small) by adding regularization with different amounts of supervised data. Additionally, we prove that decoding with the shallow fusion of external n-gram LM and automatically generated named entities always improves the performance of models, independent of the quality of pseudo-labels.
Limitations
Ethical Considerations
Acknowledgements
This work was supported by the Idiap & Uniphore collaboration project. Part of the work was also support supported by EU Horizon 2020 project ELOQUENCE [^23] https://eloquenceai.eu/ (grant number 101070558).
Appendix A Streaming decoding configurations
  • Decode chunk size = 320ms with left context of 2560ms, 5120ms and full;
  • decode chunk size = 640ms with left context of 2560ms, 5120ms and full;
Appendix B Extended results for models trained on PL data
@@
Offline models evaluation
Streaming models evaluation